Subject: 3.5-m Users Committee minutes 10/17/06
From: Bruce Gillespie
Submitted: Wed, 25 Oct 2006 15:56:24 -0600
Message number: 991
(previous: 990,
next: 992
up: Index)
APO 3.5-m Users Committee Phonecon, 10/17/06
Attending: Suzanne Hawley, Bruce Gillespie, Michael Strauss, Scott
Anderson, Jon Fulbright, John Bally, Al Harper, Jon Holtzman
Absent: Rene Walterbos, Russet McMillan, Don York, Remy Indebetouw
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Colorado (Bally) - John had no user feedback to report. He did
mention that there is a new problem with NIC-FPS Leach controller
that is affecting performance. Stephane Beland is going to diagnose
and recommend a work around until the problem can be fixed. John also
said that the NIC-FPS Fabry-Perot has been sent to IC Optical for
evaluation and possible repairs. CU is bringing a large group of
students to APO for orientation and training in a couple of weeks.
JHU (Fulbright) - Jon said that his users say things are OK, and
he'll get us a JHU science highlight for the annual report.
Princeton (Strauss) - Michael said everything is fine, and the PU
science highlight is on the way.
Chicago (Harper) - Al said that there were no new issues.
NMSU (Holtzman) - Jon mentioned that he had just had a large NMSU
student group at APO. He said that during cloudy weather he
practiced with NIC-FPS and found various performance issues [see
above] and that its on-line documentation is sparse. Suzanne
mentioned that Gabrelle Saurage (the new Obs. Specialist) is working
on instrumentation web documentation and will have new materials on
line shortly. Bruce mentioned that the observatory has not yet
officially accepted NIC-FPS because of documentation and the problem
with the Fabry-Perot.
Washington (Anderson) - Scott said he'd heard very positive feedback
from George Wallerstein about the recent excellent seeing at the
telescope.
Virginia (Indebetouw) - no report available
**********************************
Discussion of telescope/instruments report:
The report is given below; there was no discussion:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights,
9/16/06 through 10/13/06
Bruce Gillespie
0) Overview
With unflinching dedication and long hours of effort by the
engineering staff and Observing Specialists, the new top end was
successfully installed and put into service. Initial on-sky usage
subjectively suggests that the stiffness and positioning performance
of the 2ndary mirror is improved, but more testing and a longer time
baseline will be needed to characterize the improvements
definitively. The weather also cooperated as we came out of
shutdown, which allowed ample on-sky commissioning tests and
adjustments prior to returning to science observing, on schedule.
Aside from some unexpected instrument problems, the rest of the
period has been spent in summer-shutdown recovery and preparing for
the onset of winter.
1) Telescope
Aside from the new top end being put into service, the telescope
experienced some "waking-up" troubles coming out of shutdown, most
notably an altitude drive problem that was ultimately diagnosed to be
a bad DC motor, which was replaced. There is still a fair number of
little clean-up tasks left to do, and we are beginning to collect
performance data (e.g., focus with temperature, 2ndary transverse-
motion collimation [new!], wind-induced image jitter, etc.). One
observer has reported one-half arcsecond seeing at the echelle, which
may be a harbinger of good things to come.
2) Instruments
DIS spectrographs were given their seasonal refocussing, and found
not to need very much adjustment. DIS engineering time to complete
the optics and detector upgrades is scheduled for the week of 6 - 10
November.
The echelle auto-fill LN2 system became non-functional due to a
faulty sensor, allowing the dewar to warm. This necessitated
repumping the dewar before recooling. All seems OK with the system now.
SPIcam has been showing signs of cold-head temperature fluctuating.
This may be due to problems with its ion pump; diagnosis is underway,
but the instrument is otherwise usable.
The cabling for the Goddard Fabry-Perot was reworked, and the
instrument will be checked out by the PI team shortly.
3) Engineering and CIF projects
The following message about the major CIF project this year,
replacing the
top end of the telescope, was sent to apo35general on 20 Sept,
repeated here for the record:
"After years of discussion, and nearly two in design and
construction, we are pleased to announce that the APO 3.5-m telescope
has a new "top end" consisting of a new square frame, vanes, mirror
cage and actuators. The redesign greatly improves the safety of the
glass, and additionally the new top end has stiffer components and a
new transverse actuator system for the secondary mirror, both of
which promise to deliver better image quality. Also the new axial
actuators should allow smoother, and backlash free, focussing
adjustments which we have been waiting for in order to consider
implementing an autofocus capability on the telescope.
Many people have contributed to the success of this project, which
was led by Mark Klaene. We also send our thanks to Larry Carey, Jon
Davis, Dave Woods, Eric Sandberg, Russell Owen, Fritz Stauffer, Jon
Brinkmann, Ben Harris, Jim Davenport, Michael Spector, Craig Loomis,
Russet McMillan, Bill Ketzeback, Gabrelle Saurage, Jack Dembicky,
Gretchen Van Doren, Mike Evans and the staff at the UW Physics
Machine Shop for their contributions, large or small, to this fine
accomplishment.
The 3.5-m telescope will return to normal science operation on
schedule tonight. Along with the work on the new top end, numerous
other maintenance tasks were performed on the telescope and
instruments during the sumemr shutdown. We ask that observers be
alert and communicate to the Observing Specialists any unusual
behavior in the telescope or data systems, and also to exercise
patience as we perform additional necessary engineering tasks in
coming out of the shutdown.
Suzanne Hawley, APO 3.5m Director
Bruce Gillespie, APO Site Operations Manager"
4) Miscellaneous
The APOLLO program has transitioned into its near-normal operational
style of short observing runs two or three times a week. The system
is being operated by APO staff (led by Russet McMillan). The APOLLO
"Leopard" laser was given its "50,000-photon" overhaul by company
representatives.
Also, the telescope, instruments, and enclosure were given their
annual cleanings.
**********************************
Future of old top end - Hawley:
Suzanne announced that we need to make a decision about what to do
with the old top end. If we don't want to use if for a special
project, e.g., a chopping 2ndary mirror, a prime-focus camera, etc.,
we should remove it from the current location (on the secondary
exchange platform), so that we don't have a large, unnecessary
thermal mass near the telescope. Moving it from the 2ndary exchange
platform will require cutting it up, rendering it scrap. If we're
not going to use it, we should get rid of it, if only to get rid of
the unnecessary thermal mass near the telescope.
There was a limited discussion of possibilities, but no clear
winners. Suzanne said we'd wait a month, and if no compelling
suggestions come forward, we'll scrap the old top end.
**********************************
Request for 2006 science highlights - Hawley:
Suzanne reminded the group that we need one or two science highlights
from each institution for the annual report to the ARC Board of
Governors. [ed. note: as of this writing, we have highlights from
everyone except the Goddard team that uses the GFP]
**********************************
Preliminary view of next year's projects - Hawley, Gillespie:
Suzanne listed the projects being planned for next year:
Telescope
- Refurbish old drive boxes to use as spares
- Upgrade NA2 rotator (needed for TripleSpec)
- replace drive axis controllers (old ones obsolete)
- as time permits, study autofocus, and new tertiary mount
Instruments
- receive and commission TripleSpec
- receive and commission Agile (high-speed photometer)
Someone asked if TripleSpec and Agile will be TUI-compatible. The
answers are yes for TripleSpec, and maybe for Agile if it becomes
widely used.
**********************************
ARC-related news (CfP, Futures) - Hawley:
Suzanne updated the status of the post-2008 Call for Proposals review
and the Futures Committee work. The CfP review panel is close to a
recommendation, which will be sent to the Board hopefully next week.
The recommendations will be discussed at the annual Board meeting
next month. We also expect the Futures Committee report to be
finalized prior to the Board meeting.
**********************************
[open from previous months]:
ACTION: Bruce to discuss issues of computer and network security
with Fritz Stauffer. STATUS: Working group (site staff and users)
formed to frame requirements. Fair amount of work has been done, and
will be reported on at next meeting. Worth mentioning here that the
new user computer newton has been installed in the observing room
replacing tycho. At the next meeting we will discuss how users can
access it remotely
ACTION: Users Committee members should send Michael
(strauss@astro.princeton.edu) the names and e-mail addresses of 3.5-m
users who are not yet signed up for the apo35general mailer.
STAUS: Michael sent the list of mailer recipients to the committee
members for review and editing, has received feedback from a couple
of institutions.
[new actions from last month]:
ACTION: Users Committee representative to gather one or two science
highlights and send them to Bruce and Suzanne not later than mid-
October.
STATUS: Nearly closed.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 20 November, at
8:30 AM Pacific Time. Agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 991 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO